# ToddlerBot: Open-Source ML-Compatible Humanoid Platform for Loco-Manipulation

### v3

@@ Abstract
Learning-based robotics research driven by data demands a new approach to robot hardware design—one that serves as both a platform for policy execution and a tool for embodied data collection. We introduce ToddlerBot, a low-cost, open-source humanoid robot platform designed for robotics and AI research. ToddlerBot enables seamless acquisition of high-quality simulation and real-world data. The plug-and-play zero-point calibration and transferable motor system identification ensure a high-fidelity digital twin and zero-shot sim-to-real policy transfer. A user-friendly teleoperation interface streamlines real-world data collection from human demonstrations. With its data collection ability and anthropomorphic design, ToddlerBot is ideal for whole-body loco-manipulation research. Additionally, ToddlerBot’s compact size ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensures safe operation in real-world environments. Reproducibility is achieved with entirely 3D-printed, open-source design and off-the-shelf components, keeping the total cost under $6,000\leavevmode\nobreak\ \mathrm{USD}$. This allows assembly and maintenance with basic technical expertise, as validated by successful independent replications of the system. We demonstrate ToddlerBot’s capabilities through arm span, payload, endurance tests, loco-manipulation tasks, and a collaborative long-horizon scenario where two robots tidy a toy session together. By advancing ML-compatibility, capability, and reproducibility, ToddlerBot provides a robust and scalable platform for policy learning and execution in robotics research.
@@ 1 Introduction
Inspired by recent quadrupeds [1, 2] and robotic manipulators [3, 4, 5, 6, 7], ToddlerBot is designed with ML-compatibility in mind. Researchers can use ToddlerBot to not only execute policies but also to serve as a robust data collection platform. It enables the acquisition of high-quality simulation data through a plug-and-play zero-point calibration procedure and transferable motor system identification (sysID) results. These tools ensure a high-fidelity digital twin. We validate the quality of simulation data with keyframe-interpolated motions (e.g., push-ups and pull-ups) and reinforcement learning (RL) policies (e.g., walking), demonstrating the capability for zero-shot sim-to-real transfer. ToddlerBot can also acquire scalable real-world data. We design an intuitive teleoperation interface that allows simultaneous control of ToddlerBot’s upper and lower body to collect whole-body manipulation data and develop effective visuomotor policies. Additionally, ToddlerBot’s small size and weight ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensure safe and accessible operation in real-world environments.
@@
Reproducibility is achieved through low-cost, open-source designs and readily accessible hardware components. ToddlerBot uses commercially available motors and is completely 3D-printed, with a total cost under $6,000\leavevmode\nobreak\ \mathrm{USD}$ ($90\%$ of the cost is for motors and computers). We release digital twin, learning algorithms, hardware designs, and comprehensive tutorials to ensure that ToddlerBot can be built at home with basic knowledge of hardware and software, without requiring specialized equipment for manufacturing or repair. To validate reproducibility, we enlisted a CS-major student not involved in this project to independently build another instance of ToddlerBot. and successfully zero-shot transfer loco-manipulation policies between the two instances. We also received five reports of independent replications of ToddlerBot within a week by teams worldwide.
@@ 3.1 Reproducibility - A Hard Constraint
Unlike most prior works, we treat reproducibility as a hard constraint, as our humanoid platform holds no value if others cannot reproduce it. We define reproducibility as the ability of a single person to replicate the robot system at home without specialized equipment. 3D printing has become a popular method for reproducing open-source hardware systems due to its accessibility and fast turnaround time [2, 5, 4]. Despite these merits, it’s challenging to ensure sufficient strength and precision for printed plastic parts. We discuss our justifications for 3D printing in Appendix 8.1. For performance-critical components like motors and bearings, we limit ourselves to off-the-shelf parts. The total BOM cost of ToddlerBot is $6,000\leavevmode\nobreak\ \mathrm{USD}$, with 90% spent on the computer (NVIDIA Jetson) and motors (Robotis Dynamixel). We detail our design considerations in Appendix 8.2.
@@ 5 Experiments
Figure 3: Arm Span, Payload, and Trajectory Tracking. On the left, we show that with a torso dimension of $13\times 9\times 12\leavevmode\nobreak\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, about 14 times the torso size. Additionally, ToddlerBot can lift weights up to $1484\leavevmode\nobreak\ \mathrm{g}$, which is 40% of its body weight ($3484\leavevmode\nobreak\ \mathrm{g}$). On the right, we present ten consecutive real-world rollouts of an RL walking policy tracking a square trajectory with a predefined velocity profile. Both raw and smoothed linear and angular velocity tracking are displayed, with real-world results averaged across trials.
Capability: Arm Span, Payload, and Endurance. To evaluate ToddlerBot’s arm span, we teleoperate it to hug a large object using the compliant palm gripper while maintaining balance. We show that, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, approximately 14 times its torso volume.
The payload test assesses both the upper body’s lifting capacity and the lower body’s ability to maintain balance. ToddlerBot successfully lifts up to $1484\leavevmode\nobreak\ \mathrm{g}$, 40% of its total weight ($3484\leavevmode\nobreak\ \mathrm{g}$). To eliminate friction effects, we use a 3D-printed cup for the gripper to lock in securely. We incrementally add screws to the cup until ToddlerBot falls over. Results are shown on the left of Figure 3.
@@
ML-Compatibility: Omnidirectional Walking. To demonstrate ToddlerBot’s locomotion capabilities, we train RL walking policies to follow a square trajectory with a predefined velocity profile. The right part of Figure 3 presents the results, with real-world tracking data collected via motion capture. Due to RL policy limitations, both simulation and real-world tracking deviate from the command, primarily because the learned walking policy struggles with in-place rotation, causing translation offsets. However, the sim-to-real gap is notably smaller than the tracking gap, supporting a successful zero-shot transfer. Additionally, we report a position tracking error variance of $0.018\leavevmode\nobreak\ \mathrm{m}$, which demonstrates good repeatability (more details in Appendix 8.14).
ML-Compatibility: Vision-based Manipulation. We show ToddlerBot’s ability to perform bimanual manipulation and full-body manipulation by transferring octopus toys from a table and the ground to a wagon. Both tasks are trained with RGB-based diffusion policy [48] with 60 demonstrations. Across 20 test trials, we achieve a 90% success rate for bimanual manipulation and 75% for full-body manipulation. We leverage a combination of open-loop motions and closed-loop policies to enhance data collection efficiency. In the bimanual task, the torso rotating and releasing motions are open-loop, while in full-body manipulation, kneeling down is open-loop. ToddlerBot’s onboard computing runs a 300300M parameter diffusion policy with about $100\leavevmode\nobreak\ \mathrm{ms}$ latency, enabling real-time operation. Results are shown in Figure 4.

### v4

@@ Abstract
Learning-based robotics research driven by data demands a new approach to robot hardware design—one that serves as both a platform for policy execution and a tool for embodied data collection. We introduce ToddlerBot, a low-cost, open-source humanoid robot platform designed for robotics and AI research. ToddlerBot enables seamless acquisition of high-quality simulation and real-world data. The plug-and-play zero-point calibration and transferable motor system identification ensure a high-fidelity digital twin and zero-shot sim-to-real policy transfer. A user-friendly teleoperation interface streamlines real-world data collection from human demonstrations. With its data collection ability and anthropomorphic design, ToddlerBot is ideal for whole-body loco-manipulation research. Additionally, ToddlerBot’s compact size ($0.56\penalty 10000\ \mathrm{m},3.4\penalty 10000\ \mathrm{kg}$) ensures safe operation in real-world environments. Reproducibility is achieved with entirely 3D-printed, open-source design and off-the-shelf components, keeping the total cost under $6,000\penalty 10000\ \mathrm{USD}$. This allows assembly and maintenance with basic technical expertise, as validated by successful independent replications of the system. We demonstrate ToddlerBot’s capabilities through arm span, payload, endurance tests, loco-manipulation tasks, and a collaborative long-horizon scenario where two robots tidy a toy session together. By advancing ML-compatibility, capability, and reproducibility, ToddlerBot provides a robust and scalable platform for policy learning and execution in robotics research.
@@ 1 Introduction
Inspired by recent quadrupeds [1, 2] and robotic manipulators [3, 4, 5, 6, 7], ToddlerBot is designed with ML-compatibility in mind. Researchers can use ToddlerBot to not only execute policies but also to serve as a robust data collection platform. It enables the acquisition of high-quality simulation data through a plug-and-play zero-point calibration procedure and transferable motor system identification (sysID) results. These tools ensure a high-fidelity digital twin. We validate the quality of simulation data with keyframe-interpolated motions (e.g., push-ups and pull-ups) and reinforcement learning (RL) policies (e.g., walking), demonstrating the capability for zero-shot sim-to-real transfer. ToddlerBot can also acquire scalable real-world data. We design an intuitive teleoperation interface that allows simultaneous control of ToddlerBot’s upper and lower body to collect whole-body manipulation data and develop effective visuomotor policies. Additionally, ToddlerBot’s small size and weight ($0.56\penalty 10000\ \mathrm{m},3.4\penalty 10000\ \mathrm{kg}$) ensure safe and accessible operation in real-world environments.
@@
Reproducibility is achieved through low-cost, open-source designs and readily accessible hardware components. ToddlerBot uses commercially available motors and is completely 3D-printed, with a total cost under $6,000\penalty 10000\ \mathrm{USD}$ ($90\%$ of the cost is for motors and computers). We release digital twin, learning algorithms, hardware designs, and comprehensive tutorials to ensure that ToddlerBot can be built at home with basic knowledge of hardware and software, without requiring specialized equipment for manufacturing or repair. To validate reproducibility, we enlisted a CS-major student not involved in this project to independently build another instance of ToddlerBot. and successfully zero-shot transfer loco-manipulation policies between the two instances. We also received five reports of independent replications of ToddlerBot within a week by teams worldwide.
@@ 3.1 Reproducibility - A Hard Constraint
Unlike most prior works, we treat reproducibility as a hard constraint, as our humanoid platform holds no value if others cannot reproduce it. We define reproducibility as the ability of a single person to replicate the robot system at home without specialized equipment. 3D printing has become a popular method for reproducing open-source hardware systems due to its accessibility and fast turnaround time [2, 5, 4]. Despite these merits, it’s challenging to ensure sufficient strength and precision for printed plastic parts. We discuss our justifications for 3D printing in Appendix 8.1. For performance-critical components like motors and bearings, we limit ourselves to off-the-shelf parts. The total BOM cost of ToddlerBot is $6,000\penalty 10000\ \mathrm{USD}$, with 90% spent on the computer (NVIDIA Jetson) and motors (Robotis Dynamixel). We detail our design considerations in Appendix 8.2.
@@ 5 Experiments
Figure 3: Arm Span, Payload, and Trajectory Tracking. On the left, we show that with a torso dimension of $13\times 9\times 12\penalty 10000\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\penalty 10000\ \mathrm{cm}^{3}$, about 14 times the torso size. Additionally, ToddlerBot can lift weights up to $1484\penalty 10000\ \mathrm{g}$, which is 40% of its body weight ($3484\penalty 10000\ \mathrm{g}$). On the right, we present ten consecutive real-world rollouts of an RL walking policy tracking a square trajectory with a predefined velocity profile. Both raw and smoothed linear and angular velocity tracking are displayed, with real-world results averaged across trials.
Capability: Arm Span, Payload, and Endurance. To evaluate ToddlerBot’s arm span, we teleoperate it to hug a large object using the compliant palm gripper while maintaining balance. We show that, ToddlerBot can grasp objects up to $27\times 24\times 31\penalty 10000\ \mathrm{cm}^{3}$, approximately 14 times its torso volume.
The payload test assesses both the upper body’s lifting capacity and the lower body’s ability to maintain balance. ToddlerBot successfully lifts up to $1484\penalty 10000\ \mathrm{g}$, 40% of its total weight ($3484\penalty 10000\ \mathrm{g}$). To eliminate friction effects, we use a 3D-printed cup for the gripper to lock in securely. We incrementally add screws to the cup until ToddlerBot falls over. Results are shown on the left of Figure 3.
@@
ML-Compatibility: Omnidirectional Walking. To demonstrate ToddlerBot’s locomotion capabilities, we train RL walking policies to follow a square trajectory with a predefined velocity profile. The right part of Figure 3 presents the results, with real-world tracking data collected via motion capture. Due to RL policy limitations, both simulation and real-world tracking deviate from the command, primarily because the learned walking policy struggles with in-place rotation, causing translation offsets. However, the sim-to-real gap is notably smaller than the tracking gap, supporting a successful zero-shot transfer. Additionally, we report a position tracking error variance of $0.018\penalty 10000\ \mathrm{m}$, which demonstrates good repeatability (more details in Appendix 8.14).
ML-Compatibility: Vision-based Manipulation. We show ToddlerBot’s ability to perform bimanual manipulation and full-body manipulation by transferring octopus toys from a table and the ground to a wagon. Both tasks are trained with RGB-based diffusion policy [48] with 60 demonstrations. Across 20 test trials, we achieve a 90% success rate for bimanual manipulation and 75% for full-body manipulation. We leverage a combination of open-loop motions and closed-loop policies to enhance data collection efficiency. In the bimanual task, the torso rotating and releasing motions are open-loop, while in full-body manipulation, kneeling down is open-loop. ToddlerBot’s onboard computing runs a 300300M parameter diffusion policy with about $100\penalty 10000\ \mathrm{ms}$ latency, enabling real-time operation. Results are shown in Figure 4.